If a 16.5 mL sample of 1.1 M solution of each of the following acids is reacted with 0.75 M NaOH, how many milliliters of the NaOH are required for the titration? What is the total volume (in mL) of solution at the equivalence point?

​16.5 ml H2SO4 titrated with 0.75 M NaOH

Volume of NaOH (ml)   24.2 (incorrect)

Total Volume (ml)     40.7 (incorrect)
Answer #1

     H2SO4 + 2NaOH ----------------> Na2SO4 + H2O

   1 mole       2 moles

H2SO4                                                                      NaOH

M1 = 1.1M                                                                 M2 = 0.75M

V1    = 16.5ml                                                             V2    =

n1   = 1                                                                       n2 = 2

               M1V1/n1     =        M2V2/n2

                          V2     = M1V1n2/M2n1

                                   = 1.1*16.5*2/0.75*1    = 48.4ml

volume of NaOH    = 48.4ml

total volume          = 16.5 +48.4   = 64.9ml